OpenStudy (anonymous):

find the value of x. (2x+10), 148, (2x), 112 ?

jhonyy9 (jhonyy9):

2x+10 ,148,2x,112 - 148-112=36 but because from 112 till 148 there are two steps so 36 need dividing by 2 what will be equal 18 - this mean that 2x=112+18=130 --- from this result that x=130/2 = 65 x=65 - hope you understand it